# MapFormer: Self-Supervised Learning of Cognitive Maps with Input-Dependent Positional Embeddings

### v2

@@ 1 Introduction
Figure 1: Cognitive Maps: Disentangling Structure and Content for Path Integration in Episodic and Working-Memory Models. (a) Cognitive maps model the relationships between entities, such as objects in space or people working in an organization. In a cognitive map, both the entities and their position are represented but they can be factorized from each other. The position (or, more generally, the ‘state’) in a cognitive map is updated via actions (blue arrows). For example, actions can update position on a grid (e.g., ‘step right’ in the left drawings), or in more abstract environments, on an organization chart (e.g. ‘the boss of’, on the right). Actions are assumed to have the same effect on the state in the map across all positions and across all environments (up vs bottom). Path integration can be achieved by sequentially combining actions together $a_{0\to T}=(a_{0}\to...\to a_{T})$ (green arrow). (b) (Top) The mental trajectory of a rodent navigating a maze lies on a torus (figure adapted from Gardner2022 under $C$ BY 4.0 license); (bottom) trajectories during a 2D navigation task in MapFormers also lie on a torus. The red curve represents a square trajectory performed by the simulated agent during a navigation task, returning to its initial starting point, as desired. (c-d) An illustration of the conceptual framework behind MapFormers, and of the differences between Episodic Memory (EM) and Working Memory (WM) models Whittington2025. (c) Episodic Memory (EM) model: Positions (or ‘states’, more generally) in a cognitive map $p_{t}$’s (in blue), at times $t$’s, are encoded in a separate neural space by a dedicated neural population, which does not directly interact with the input $x_{t}$ (red), being updated via actions only. Actions $a_{t}$’s, when they occur in the input, can update the position via input-dependent matrices $W_{a_{t}}$ (green): $p_{t}=W_{a_{t}}p_{t-1}$. $W_{a}$’s are input-dependent matrices and are computed through an exponential map, which maps between abstract representations of actions $A_{t}:=\Delta_{t}\omega S=\theta_{t}S$ (in the Lie-Algebra) and the corresponding weight matrix $W_{a_{t}}:=\exp(A_{t})=R_{\theta_{t}}$ (in the corresponding Lie group; section 3.3). The exponential mapping is computed based on the integration time $\Delta_{t}$ of the action (light green) – roughly, how long the synaptic weights $S$ need to be applied. Finally, the conjunction (purple) of the content (what) and the structure (where) can be computed. (d) Working Memory (WM) model: In contrast to the EM model, which factorizes content (red) and position (blue) in two separate neural spaces, MapFormers-WM represent position implicitly. Actions in the WM model directly update the conjunction of what and where (purple). The conjunction states $h_{t}$’s in the WM model are thus smaller by a quadratic factor compared to the EM model.
The ability of Transformers NIPS2017_3f5ee243 to process sequences in parallel has enabled the training of Large Language Models (LLMs) on increasingly large datasets, creating models able to write poetry, solve complex mathematical problems, or even describe what they observe in an image, results so impressive that debates about their reaching AGI (Artificial General Intelligence) have already started bubeck2023sparksartificialgeneralintelligence; morris2024levelsagioperationalizingprogress. Yet, the same models learn incoherent world models vafa2024evaluatingworldmodelimplicit, make logical mistakes on tasks that are trivial to us humans williams2024easyproblemsllmswrong, and even when they are right, they can be so for the wrong reasons mccoy-etal-2019-right. Moreover, the fact that Transformers fail to generalize OOD to non-regular formal tasks deletang2023neural shows that some key structural biases are most likely missing for transformer-based models to reach human-like generalization abilities and AGI.
In comparison, human and animal cognitive abilities appear much more robust and flexible. It has been suggested that the reason for this is our ability to build cognitive maps BEHRENS2018490, internal models of the relationships among entities or events in the world (Figure 1a). This notion was first introduced in 1948 by Edward Tolman Tolman1948-TOLCMI as a way to explain how rats could successfully find new routes in a maze after old ones were blocked, behavior that cannot be explained by simple conditioning, suggesting that rats develop an internal cognitive map of their environment to flexibly adapt to new scenarios. Neural evidence of internal maps was first discovered in the hippocampus O'Keefe1978, with place cells firing every time a rodent came back to a specific location. Later, grid cells were found in the entorhinal cortex (EC) Hafting2005, firing periodically with a “grid-like” pattern as an animal was exploring a room, adding further evidence of an internal map within mammal’s brains.
Cognitive maps could be important for generalization and flexible decision-making, as they provide an internal model that represents the structural relationships among entities in the world. This abstraction holds true for both spatial navigation (e.g., finding a shortcut to get back home, or finding a detour around an obstacle) and non-spatial, conceptual information (e.g., inferring family relations). For example, in the case of spatial navigation, if a wall blocks the road ahead, understanding that the sequences $(\rightarrow,\rightarrow)$ and $(\uparrow,\rightarrow,\rightarrow,\downarrow)$ are equivalent allows an agent to ignore unexpected obstacles by simply walking around before continuing on the desired path. If someone says “Bob is the brother of Alice’s mother,” when asked “Who is Alice’s uncle?” one can easily reply “Bob” by simply knowing the abstract relationships of a family tree. Seeing structure as a connected graph of relations unifies spatial and non-spatial cognitive maps, and is the reason why some have theorized that the neural mechanism allowing the creation of spatial cognitive maps might be “recycled” in humans to build abstract knowledge graphs Aronov2017; KRIEGESKORTE2016280; Park2021 (fig. 1(a)).
Cognitive maps achieve robust generalization by separating the abstract, invariant structure of the world from the transient, variable content of observations BEHRENS2018490. Because cognitive maps define the position of an object via its relations to other entities in the world, these models can be formalized as graphs, whose nodes represent objects and edges the relations connecting them Whittington770495. This factorizes what has been identified from where it was seen. Making structure independent from content dramatically helps generalization, assuming that the structural relations learned in one environment remain true in new ones. Indeed, making a step right then left brings us to the same point whether we perform it in Paris or New-York, while our representation of numbers remains the same whether we are counting apples or oranges. Disentangling structure from content also decreases the required amount of training data by a quadratic factor, since the joint distribution can be expressed as the product of its two marginals: $P(x_{i},p_{j})=P_{x}(x_{i})P_{p}(p_{j})$. In a world where structure remains stable but observations are infinite, this factorization is fundamental.
In addition to learning the structural relations among entities, an agent must also learn the set of actions that affect its state within that structure and how these actions combine. Learning the effect of actions on the state is important for efficient encoding of the underlying structure. For example, knowing how the four cardinal directions (North, South, East, West) combine is enough to define any new trajectory on a grid, while the group of integers $(\mathbb{Z},+)$ needs only 0, Peano’s successor operation ++ and its inverse - to infer an infinity of numbers. Once structural relations and actions are represented, an agent can use them to mentally keep track of its position (or more generally, the state) and to simulate the impact of future trajectories olaf2018, a process known as path integration (fig. 1(a)). Humans and animals are highly proficient at path integration Hafting2005; HOWARD20141331; Wehner1981, which provides two key advantages. First, it allows the agent to maintain an accurate internal estimate of its location even without external sensory input. Second, this ability to mentally integrate actions along the structure allows for the representation of planned actions, which is essential for flexible decision-making and rapid adaptation to unexpected obstacles.
@@
What kind of structural bias must be introduced into a deep learning architecture to enable it to effectively learn a cognitive map and perform robust path integration? Previous work made progress in modeling cognitive maps with RNNs with action-dependent matrices Whittington770495, allowing them to solve structure-dependent tasks and generalize to unseen scenarios. This work also explains the duality between episodic and working memory (EM / WM) Whittington2025: EM uses action-dependent matrices to sequentially update a separate pool of neurons representing the agent’s abstract location in a cognitive map, used for long term recall of structured memories stored in synaptic weights (fig.1(c)); WM uses these matrices to directly move the state into distinct neural subspaces — acting as memory slots — allowing the model to keep the whole structure in neural activity (fig.1(d)). Although very promising, these models do not scale since (1) their action-dependent matrices need to be sequentially multiplied to perform path integration, making them slow to train and incompatible with parallel sequence processing, and (2) the models can only be applied to problems where the underlying structure is already known (for example, where the number of actions is known rather than learned).
@@
To address this, we introduce MapFormers, a class of transformers that learn context-dependent cognitive maps with a generic neural mechanism, allowing the learning of structure at larger scale. MapFormers compute path integration in parallel and therefore allows scaling up to complex problems. Furthermore, MapFormers are trained in a fully self-supervised manner, and learn by themselves to disambiguate between actions and observations, which have opposite effects on the learned representations. The former update the internal state’s structure via learned operations but leave its semantic content untouched, while the latter leave the internal structure fixed but change the state’s content. Empirically, we argue that MapFormers, to the best of our knowledge, are the first generic architectures to learn cognitive maps and generalize to longer sequences without supervision, and with minimal assumptions about the underlying structure. We introduce two variants of MapFormers, which unify relative rotary (RoPE su2023roformerenhancedtransformerrotary) and absolute positional embeddings, respectively as models of working and episodic memory, confirming the fundamental role that action-dependent positional embeddings play in encoding the location of tokens within a context-dependent cognitive map whittington2022relatingtransformersmodelsneural. Finally, we draw theoretical arguments from Lie-group theory to characterize the class of structures that sequence-processing models are able to learn, allowing us to show that diagonal matrices used in current selective State Space Models are not expressive enough to learn cognitive maps.
@@ 2 Related Literature
In a seminal work Whittington770495, the authors presented a set of tasks, for both spatial (navigation) and non-spatial (e.g., family trees) problems, which require learning a cognitive map in order to achieve strong generalization abilities. The Tolman-Eichenbaum Machine (TEM) was suggested as a model that can learn cognitive maps to solve these tasks. Specifically, TEM is an RNN-based architecture that sequentially updates its internal position with action-dependent weights, before querying a Hopfield memory module ramsauer2021hopfieldnetworksneed with its updated position to retrieve what has been seen at a previously visited location.
To extract and build a cognitive map from observational input and, in particular, to learn how an action taken affects the current position in the map, TEM suggests moving beyond standard RNNs. In standard RNNs, the recurrent weight matrix is fixed and applied to any input in the same way. However, rules need to be dynamically applied or not depending on context. In order to model how input actions update the current position in a cognitive model, the weight matrix must depend on the input. That is, instead of having a fixed weight matrix $W$, the model needs to learn an input dependent matrix, specific to each action.
The TEM framework was originally suggested as a computational model of the entorhinal and hippocampus regions in the Medial Temporal Lobe (MTL) for the encoding of cognitive maps in long-term episodic memory (EM), explaining the role of the hippocampus in spatial and non-spatial generalizations. In a subsequent extension of this work, Whittington2025 broadened the framework to explain how cognitive maps can also be acquired by a different system, proposing a model of the role of the prefrontal cortex (PFC) for working memory (WM) for sequence memory tasks. The WM-based model achieves this without an explicit separation of the content from the learned map structure, instead employing dynamic encoding mechanisms to update structured activity slots.
However, a key limitation of models based on standard RNNs — including both the EM and WM architectures — for large-scale applications, is the inherent computational bottleneck created by their sequential processing. This is particularly restrictive in real-world scenarios, such as long-range path integration, where the cumulative effect of a sequence of actions requires the slow, iterative multiplication of recurrent weight matrices. This is true also for more recent work using Transformers - In a follow-up work, whittington2022relatingtransformersmodelsneural introduced TEM-t, a model for TEM that can be seen as a Transformer, where positional embeddings in the model represent abstract positions in a cognitive map. However, TEM-t cannot leverage parallel sequence processing abilities of Transformers since it requires sequential updates of positional embeddings with input-dependent matrices. MapFormers overcome this limitation by employing cumulative summation across the entire sequence, allowing the computation of path integration to occur in parallel, as we will see below.
Positional embeddings were traditionally added to Transformers because the attention mechanism is permutation invariant, with position either added explicitly within a separate pool of neurons NIPS2017_3f5ee243, or implicitly by representing the relative distance between keys and queries su2023roformerenhancedtransformerrotary; shaw2018selfattentionrelativepositionrepresentations. Positional embeddings factorize content from position in transformers, but since these positional representations are shared by all sequences, and are not input-dependent, they cannot represent the task’s latent structure. In line with TEM-t, other work has demonstrated the generalization power of sequence-dependent positional encoding. Specifically, Contextualized equivariant Position Encoding (TAPE) zhu2025rethinkingaddressinglanguagemodels introduced input-dependent MLP after the attention mechanism, updating positional embeddings for the next layer, which reverses the logic of TEM-t, where positions are modified before querying the attention module. CoPE golovneva2024contextualpositionencodinglearning modifies its positional encoding by compressing the relative distances between keys and queries, achieving almost perfect length generalization in some 1-dimensional tasks. PoPE gopalakrishnan2025decouplingwhatwherepolar follows a logic similar to CoPE, by directly modifying RoPE’s relative distances, meaning that in both CoPE and PoPE, relative distances are not constrained to be integers and can take any positive real value, allowing the model to shorten the distance between structurally similar tokens. While these three methods (TAPE, CoPE and PoPE) yield superior results in length generalization and a selection of formal tasks, their representation power is poorly understood. Specifically, little is known about what kind of structural features these model can or cannot learn. Also, no connections have been made to cognitive maps and path integration, which is one of the main goals of MapFormers.
Figure 2: MapFormers: Path-Integration Transformer-Based Models that Learn Cognitive Maps. Overview of our Transformer architectures unifying working and episodic memory as relative and absolute positional encodings, respectively. In both case, rotation angle $\theta$ is obtained via a low-rank projection of input $X$, before applying a cumulative sum (cumsum) along the temporal dimension, to perform path integration (PI) in parallel $\theta^{PI}=\omega\Delta^{PI}=\omega\sum_{s\leq T}\Delta_{s}$. Input-dependent block-diagonal matrices computed via $R_{\theta^{PI}}=\exp({\theta^{PI}}S)$ ($S:=-S^{\top}$) capture the task’s latent structure, while diagonal matrices used in selective SSMs do not. (a) Working Memory MapFormer (MapWM) with relative positional encoding, where $Q,K$ are rotated via input-dependent matrix $R_{\theta_{PI}}$. In MapFormers, $\theta$ is learned, but keeping it fixed in MapWM gives exactly RoPE (b) Episodic Memory MapFormer (MapEM) with two parallel attentions $A_{X}$ and $A_{P}$. The former captures raw similarities between external inputs and the latter encodes the problem’s structure. Combining these two attention matrices with element-wise multiplication $A:=A_{X}\odot A_{P}$ is equivalent to computing a single attention $A_{G}$ where keys and queries are the conjunction of observation and position: $g_{t}=\text{vec}(x_{t}^{\top}.p_{t})\in\mathbb{R}^{d^{2}}$, like in TEM-t whittington2022relatingtransformersmodelsneural. In MapEM, absolute position embeddings is obtained via path integration of initial and shared key/query positions $Q_{P}^{*}$ and $K_{P}^{*}$.
@@ 3 General Setup
This section provides the theoretical details that motivate the development of MapFormers, as well as their implementation details. Section 3.1 describes the tasks we used for evaluating the models: (1) A selective-copy task and (2) Forced-Navigation tasks. Section 3.2 describes the framework for action-dependent matrices and their formulation as a group. Section 3.3 introduces the idea that for learning a cognitive map, path integeration can be more efficiently computed at the Lie algebra vector space, rather than at the group level. Finally, given these insights from Lie theory, 3.4&3.5 show how to build state-space models (SSMs) and Transformer-based models, respectively, that can learn a cognitive map.
3.1 Task
Selective Copy
@@ 3.1 Task
This task was introduced in Mamba gu2024mambalineartimesequencemodeling and later used in CoPE golovneva2024contextualpositionencodinglearning to ensure that models are able to ignore distractors, which transformers with static positional encoding like RoPE su2023roformerenhancedtransformerrotary are unable to do. The model receives a sequence of symbols and must copy it entirely while ignoring a blank token $B:CFEABABBBCBF\to CFEAACF$. Intuitively, the model has to "compress" the distances between tokens, and increment its position within the cognitive map (here the number line) every time it sees a non blank token (incrementing the number of tokens to copy by one). Solving this task requires a flexible gating mechanism in positional encoding that baseline transformers do not possess.
Forced-Navigation Task
Following the task proposed by Whittington770495, we formulate the learning problem within a standard self-supervised setup. An agent is presented with a set of inputs: $s_{1},s_{2},\dotsc$ and its goal is to predict the next input token. We assume that this stream of inputs can contain both sensory observations (e.g., a dog) and actions (e.g., move left). To efficiently solve the task, and generalize OOD, the model needs to learn the hidden structure in the input (e.g., the position of the dog and the relationship of this position to other positions of other input tokens) and separate it from its contingent content (i.e., the dog itself in this position, which might be another animal in another environment). In other words, the model needs to learn to build a cognitive map. Learning this cognitive map also means to identify which input tokens $s_{t}$ correspond to actions, as well as learning how these actions affect the current, ‘active’, position in the cognitive map.
It is important to note here that even though the model receives a sequence of symbols and effectively does not purposely take any action, it must nonetheless infer the meaning of each symbol without any supervision. Indeed, it sometimes happen to simply follow a schematic sequence of instructions (ex: "Turn right then left at the end of the tunnel") and infer one’s position accordingly.
3.2 Action-Dependent Matrices Form a Lie Group
As discussed above, the key idea to learn a cognitive map is to move beyond fixed matrices and learn action-dependent matrices, specific to each input action. In dorrell2023actionableneuralrepresentationsgrid, the authors introduce actionable representations using group and representation theory, as a formalism to map an input action $a\in\mathbb{R}^{r}$ (e.g. in 2D, $r=2$ and ‘move right’) to a neural transformation $W_{a}\in M_{n}(\mathbb{R})$ (e.g., the corresponding weight matrix of a neural network that implements the effect of this action on the represented position in space $p$):
$\phi:\mathbb{R}^{r}$ $\to M_{n}(\mathbb{R})$ (1)
$a$ $\mapsto W_{a}$
The insight that action-dependent matrices need to form a group, and in particular a Lie group over $\mathbb{R}$, allows the use of results from group and representation theory to understand what are the possible constraints on the form of $W_{a}$’s dorrell2023actionableneuralrepresentationsgrid. Specifically, the Peter–Weyl theorem knapp1996lie states that any compact group $\mathbf{G}$ can be represented in $M_{n}(\mathbb{R})$ as:
where $\omega_{k},\Delta_{t_{k}}\in\mathbb{R}^{r},r=1,2$ are the rotation durations and frequencies, respectively, the latter representing movement at different scales. That is, in this example, each $I_{k}$ performs a rotation of a subspace of the neural space. The matrix $M$ then performs a change of basis on this result, assuming $M$ is not the identity.
3.3 Path Integration Can Be More Easily Computed in the Lie-Algebra Space
To learn a cognitive map, we saw that it is desirable to require action-dependent matrices to form a Lie group $\mathbf{G}$. A Lie group is a group $G$ that is also a smooth manifold, such that the group operations (multiplication and inversion) are smooth maps. For example, the group of 2D rotations $\mathbf{SO}(2)$ is a Lie group. To build a neural model that can learn cognitive maps, the goal would be to learn action-dependent matrices that form a Lie group.
To further characterize the group of action-dependent matrices, it is often simpler to analyze the corresponding Lie algebra. The associated Lie algebra $\mathfrak{g}$ of a Lie group $\mathbf{G}$ is the tangent vector space of $\mathbf{G}$ at the identity element. Lie-group theory shows the existence of an exponential map between the Lie algebra and the group $\exp(\cdot):\mathfrak{g}\to\mathbf{G}$, and in the case of a compact group, where the exponential map is surjective, any element $W_{a}\in\mathbf{G}$ of the group can be expressed as the exponential of an algebra element $A$, $W_{a}=\exp(A)\in\mathbf{G}$. This is useful since studying the structure of Lie groups (which involves, e.g., matrix multiplication) is often simpler from the point of view of the corresponding Lie algebra (which is linear and involves addition operations).
For the case of cognitive maps, we will think of elements of the Lie algebra $A$’s as action representations, and the corresponding group elements (expressed via the exponential map) as the action-dependent matrices $W_{a}$’s, which implement these actions in synaptic weights. The exponential map provides the link between the vector space of actions (the Lie algebra) and the space of weight matrices (the Lie group). Specifically, given the property of the exponential map (if the $A_{i}$’s commute): $\exp\left(\sum_{i}A_{i}\right)=\prod_{i}\exp(A_{i})$, the exponential mapping translates the additive structure of the Lie algebra (cumulation of actions in this vector space - e.g, ‘move right’ then ‘move up’) into the multiplicative structure of the Lie group (multiplication of the corresponding action-dependent weight matrices). This is useful for path integration, since the cumulative effect of taking actions $A_{1},...,A_{n}$ can be translated to the final action-dependent weight matrix $W_{a}=\prod_{i=1}^{n}\exp(A_{i})$. As we will see, the exponential map allows in some cases to compute path integration in parallel, in the vector space of the Lie algebra, by first computing the cumulative summation of subsequent actions and only then computing the exponent of the result.
For example, for spatial problems, such as navigation on a circle or a torus, the Lie group of interest is $SO(2)$ - the group of all 2D-rotations.

### v3

# MapFormer: Self-Supervised Learning of Cognitive Maps with input-dependent positional encoding

@@ 1 Introduction
Figure 1: Cognitive Maps and Path Integration: Factorizing Structure and Content. Cognitive maps factorize entities from their position (or ”state”) within a relational structure. Actions (blue arrows) update a position on the map—such as ”step right” on a grid or ”manager of” in an organization chart—and have a consistent, position-invariant effect across all environments. Path integration can be achieved by sequentially combining actions together $a_{0\to t}=(a_{t-1}\circ...\circ a_{0})$ (green arrow).
Figure 2: Disentangling Structure and Content for Path Integration in Episodic and Working Memory Transformers. Overview of our MapFormers architectures inspired by models of Episodic (EM; panel a) and Working Memory (WM; panel b) from (Whittington et al., 2025). In both cases, rotation angle $\theta=\omega\Delta$ is obtained via a low-rank projection of input $X$, before computing a cumulative sum (cumsum) along the temporal dimension, to perform path integration (PI) in parallel: $\theta^{PI}=\omega\Delta^{PI}=\omega\cdot\mathrm{cumsum}_{t}(\Delta)$. Input-dependent, block-diagonal, rotation matrices $R_{\theta^{PI}}$ represent path-integration in the learned cognitive map. (a) Episodic Memory MapFormer (MapEM) with two parallel attentions $A_{X}$ and $A_{P}$. The former captures raw similarities between external inputs and the latter encodes the problem’s structure. In MapEM, absolute position embeddings is obtained via path integration of initial position $P^{*}$, which can be initialized separately for keys and queries $Q^{*}_{P}$ and $K^{*}_{P}$. (b) Working Memory MapFormer (MapWM) with relative positional encoding, where $Q,K$ are rotated via input-dependent matrix $R_{\theta^{PI}}$. In MapWM, $\theta$ is learned, but keeping it fixed gives exactly RoPE. Diagrams created with Draw.io.
Transformer-based models have revolutionized AI, sparking debate over whether we are on the verge of achieving Artificial General Intelligence (Bubeck et al., 2023; Morris et al., 2024). Yet, beneath their impressive performance, these models remain surprisingly brittle; when carefully tested, they often fail to generalize to out-of-distribution tasks that humans and animals can handle with ease (Nezhurina et al., 2024; Zhang et al., 2024). This gap suggests that Transformers might still lack the necessary structural biases that allow biological brains to navigate unknown scenarios and reason through novel problems efficiently.
A potential explanation for this gap lies in the ability to construct cognitive maps (Behrens et al., 2018): internal models of the world that capture the relationships among entities or events in the world. Cognitive maps separate the abstract, invariant structure of an environment from its specific, variable content. By disentangling the ”where” from the ”what”, biological systems can apply a known structure to entirely new observations. Structure-content separation is important for both spatial navigation (e.g., finding a shortcut to get back home or finding a detour around an obstacle) and non-spatial, conceptual information (e.g., inferring family relations). Seeing structure as a connected graph of relations unifies such different domains: nodes represent states, while edges represent the relationships between them, spatial or conceptual. In a world where possible observations are unbounded but the underlying structure remains fixed, this factorization of structure and content provides the basis for flexible intelligence that modern AI has yet to master.
To navigate a structure, an agent must learn the set of actions that affect its state and, crucially, how these actions combine. This is called path integration (fig. 1, green arrows) – the process of integrating a sequence of actions, which is essential for tracking state changes and for mental planning (Ólafsdóttir et al., 2018). For example, knowing how the four cardinal directions (North, South, East, West) interact is sufficient to define any novel trajectory on a grid. Path integration can also enable an agent to plan shortcuts or bypass obstacles by recognizing that distinct action sequences, such as $(\uparrow,\rightarrow,\rightarrow,\downarrow)$ and $(\rightarrow,\rightarrow)$, are functionally equivalent. Similarly, in a conceptual map, one can identify ”Bob” as Alice’s uncle when told ”Bob is the brother of Alice’s mother” by mentally composing these individual social relations into a direct relational link. Humans and animals are highly proficient at this type of relational composition (Hafting et al., 2005; Howard et al., 2014; Wehner and Srinivasan, 1981), a capability for robust out-of-distribution generalization that current AI systems still struggle to replicate (Vafa et al., 2024; Kaesberg et al., 2025).
@@
What kind of structural bias must be introduced into a neural architecture to enable it to effectively learn a cognitive map and perform robust path integration? Previous work made progress in modeling cognitive maps using RNNs with action-dependent matrices (Whittington et al., 2019, 2025), allowing them to solve structure-dependent tasks and generalize to unseen scenarios. Although very promising, these models do not scale since (1) their action-dependent matrices need to be sequentially multiplied to perform path integration, making them slow to train and incompatible with parallel sequence processing, and (2) the models can only be applied to problems where the underlying structure is already known (for example, where the number of actions is known rather than learned).
@@
To address this, we introduce MapFormers, a class of transformers that learn context-dependent cognitive maps with a generic neural mechanism, allowing the learning of structure at larger scale. MapFormers compute path integration in parallel and therefore allow scaling up to complex problems. Furthermore, MapFormers are trained in a fully self-supervised manner, and learn by themselves to disambiguate between actions and observations, which have opposite effects on the learned representations. Actions update the internal state’s structure via learned operations but leave its semantic content untouched, while observations leave the internal structure fixed but change the state’s content. Empirically, we argue that MapFormers, to the best of our knowledge, are the first generic architectures to learn cognitive maps and generalize to longer sequences without supervision, and with minimal assumptions about the underlying structure. We introduce two variants of MapFormers, which unify relative rotary (RoPE; Su et al. 2023) and absolute positional embeddings, respectively as models of working and episodic memory, confirming the fundamental role that action-dependent positional embeddings play in encoding the location of tokens within a context-dependent cognitive map (Whittington et al., 2022). Finally, we draw theoretical arguments from Lie-group theory to characterize the class of structures that sequence-processing models are able to learn, allowing us to show, for example, that diagonal matrices used in current selective State Space Models are not expressive enough to learn cognitive maps.
@@ 2 Related Literature
In seminal work, Whittington et al. (2019) presented a set of tasks, for both spatial (navigation) and non-spatial (e.g., family trees) problems, which require learning a cognitive map in order to achieve strong generalization abilities. The Tolman-Eichenbaum Machine (TEM) was suggested as a model of the entorhinal and hippocampus regions, which can learn cognitive maps to solve these tasks. Specifically, TEM is an RNN-based architecture that uses action-dependent weights to update its position, before querying a Hopfield memory module (Ramsauer et al., 2021) with its updated position to retrieve what was seen at a previously visited location.
Whittington et al. (2025) extended the framework to model the role of the prefrontal cortex (PFC) in working memory, representing structure implicitly, by rotating the full neural activity into structured activity slots with action-dependent matrices. Whittington et al. (2022) show that TEM can be seen as a Transformer (TEM-t), where positional embeddings represent abstract positions in a cognitive map. However, TEM-t cannot leverage the parallel sequence processing abilities of Transformers since it requires the sequential updates of positional embeddings with action-dependent matrices, showcasing a core limitation of actionable models.
In line with TEM-t, other work has demonstrated the generalization power of sequence-dependent positional encoding in transformers (Golovneva et al., 2024; Zhu et al., 2025; Gopalakrishnan et al., 2025), yielding superior results in length generalization and a selection of formal tasks, but their representation power is poorly understood. Specifically, little is known about what kind of structural features these models can or cannot learn. Moreover, no connections have been made to cognitive maps and path integration, which is one of the main goals of MapFormers.
@@ 3 General Setup
3.1 Tasks
To study the performance and OOD generalization of MapFormers, compared to baseline models, we focused on two main tasks:
@@ 3.1 Tasks
Selective Copy This task was introduced in Mamba (Gu and Dao, 2024) and used in CoPE (Golovneva et al., 2024) to study whether the model is able to ignore distractors, which transformers with static positional encoding like RoPE (Su et al., 2023) are unable to do. The model receives a sequence of symbols and needs to copy it fully while ignoring a blank token $B:CFEABBCBF\to CFEACF$. Solving this task requires a flexible gating mechanism that baseline transformers do not possess.
Forced-Navigation Task Following the task proposed by Whittington et al. (2019), we formulate the learning problem within a standard self-supervised setup. This task consists of a random sequence of symbols $s=\left(a_{1},o_{1},\dotsc,a_{T},o_{T}\right)$ with action $a_{t}$ sampled from $\left\{\uparrow,\downarrow,\rightarrow,\leftarrow\right\}$ and observation $o_{t}$ from $\{o_{1},\dotsc,o_{K}\}$, a finite set of $K$ objects, plus a blank token $B$ when the location is empty. The model is asked to predict the upcoming observation each time it comes back to a previously visited location. To solve the task and generalize OOD, the model needs to represent the underlying 2D grid structure of the problem, independently of the specific observations (content) that appear randomly in the training data. The model also needs to understand that only action tokens update position, and distinguish them from observation tokens.
3.2 Cognitive Maps and Additive Path-Integration
We define a cognitive map as a discrete multirelational graph $G=(N,E,\mathcal{R},L)$, where nodes $N$ represent states and $E$ is the set of edges. The set $\mathcal{R}=\{r_{1},\dotsc,r_{m}\}$ denotes a finite set of cardinal relations, and $L:E\to\mathcal{R}$ is a labeling function that assigns each edge to a specific relation type. An agent navigating this map occupies a position $p\in N$ and possesses a set of cardinal actions $\mathcal{A}=\{a_{1},\dots,a_{m}\}$. Each action $a_{i}$ is defined as a mapping $a:N\to N$, such that applying an action $a_{t}\in\mathcal{A}$ at a position $p_{t}$ at time $t$ results in a new position $p_{t+1}=a_{t}(p_{t})\in N$.
In order to successfully navigate the map, the agent needs to learn the actions in $\mathcal{A}$ such that they align with the structural relations in $R$. Specifically, for any relation $r_{j\to i}\in R$ connecting $n_{j}$ to $n_{i}$, the agent needs to learn an action $a_{t}\in\mathcal{A}$ that satisfies the condition $p_{t+1}=a_{t}(p_{t}=n_{j})=n_{i}$. This alignment ensures that the agent’s internal operators accurately reflect the graph’s topology, allowing for predictable transitions through the state space.
Given $\mathcal{A}$, the agent can perform path integration (fig. 1, green arrows) by composing a sequence of actions. The resulting position $p_{i}$ reached from an initial position $p_{j}$ through a sequence of actions is defined by:
$p_{i}=a_{j\to i}p_{j}=(a_{i-1}\circ\dotsb\circ a_{j})p_{j}$
Where we denoted integrated actions as $a_{j\to i}=a_{i-1}\circ\dotsb\circ a_{j}$. Note that position is always defined by integrated action $a_{j\to i}$, representing either relative or, when referenced to the origin $p_{0}$ ($j=0$), absolute position.
In previous work, Whittington et al. (2019, 2022, 2025) modeled cognitive maps in RNNs with action-dependent recurrent weights, to compute integrated actions $a_{j\to i}$ via matrix multiplication: $\mathbf{W_{j\to i}}:=\prod_{s=j}^{i-1}W_{a_{s}}$. These matrices are required to form a Lie group, in order to have properties such as closure, an identity (null action) and inverse element, which allows—for compact groups—to represent $W_{a}$’s as block-diagonal matrices (Dorrell et al., 2023).
We pursue this line of work but avoid sequential matrix multiplications $\prod_{s=j}^{i-1}W_{a_{s}}$, which are computationally demanding, and instead compose actions linearly, allowing efficient path integration. To achieve this, we compute path integration in the Lie algebra and map the result onto the Lie group using the exponential map (sec. A.1).